### v2

# Asymptotically Optimal Lazy Lifelong Sampling-based Algorithm for Efficient Motion Planning in Dynamic Environments

Abstract
The paper introduces an asymptotically optimal lifelong sampling-based path planning algorithm that combines the merits of lifelong planning algorithms and lazy search algorithms for rapid replanning in dynamic environments where edge evaluation is expensive. By evaluating only sub-path candidates for the optimal solution, the algorithm saves considerable evaluation time and thereby reduces the overall planning cost. It employs a novel informed rewiring cascade to efficiently repair the search tree when the underlying search graph changes after evaluation. Simulation results demonstrate that the algorithm outperforms various state-of-the-art sampling-based planners in addressing both static and dynamic motion planning problems.
Index Terms:
Motion planning, lifelong planning, dynamic environments, sampling-based, asymptotically optimal, lazy search
I Introduction
Real-time motion planning is crucial for robots to navigate safely in partially known or completely unknown environments. In such scenarios, robots must interleave planning with action executions to ensure that their plans remain viable and safe, accounting for dynamic obstacles. Efficient planning algorithms are necessary in these cases. These algorithms should be capable of updating previous solutions based on new environment information in a short time.
Early researches proposed that reusing previous search can significantly enhance planning efficiency [1, 2]. The strategy, commonly known as lifelong planning in literature, was combined with graph-based path planning techniques, such as A [3]. In particular, Lifelong Planning A (LPA) [4] and its variants D/D Lite [5], repair the same search graph that approximates the problem domain throughout the entire navigation process. Any inconsistency resulting from edge changes in the graph is identified, and the inconsistencies are efficiently propagated to the relevant parts of the search graph to repair an A-like solution. It ensures efficient adaptation to the dynamic nature of the problem domain. However, LPA is only resolution optimal and complete with respect to the particular underlying graph used. In other words, they almost surely find sub-optimal paths with respect to the robot and environment. Improving completeness/optimality requires recomputing the search graph, which is infeasible in applications.
Sampling-based planning methods, such as RRT [6], RRT [7], Informed RRT [8], in contrast, avoid an explicit offline construction of the problem domain and refines an asymptotically dense Random Rapidly-explored Graph (RRG) directly in the robot’s state-space. These methods aim to find asymptotically optimal paths and probabilistically complete in terms of the robot and its environment. However, a significant drawback of these methods is their inability to efficiently handle graph changes, as they require replanning from scratch whenever a change is detected. To address this limitation, previous works have explored reusing previous search trees to improve the replanning efficiency of sampling-based planning methods [9, 10, 11, 12, 13]. Unfortunately, these methods do not guarantee the quality of their solutions. In contrast, asymptotically optimal lifelong sampling-based methods [14, 15] stand out from previous approaches by offering asymptotically optimal solutions. They use a rewiring cascade to remodel the search tree around dynamic obstacles promptly. These methods offer significant benefits, particularly in applications that require high solution quality.
Unfortunately, the existing asymptotically optimal lifelong sampling-based methods primarily focused on enhancing solution quality but need to be more indifferent to the number of edge evaluations. Specifically, they must iteratively evaluate all edges to check their feasibility before rewiring, irrespective of their relevance to the solution. In motion planning problems, edge evaluation typically consists of computationally-cost procedures, such as solving two-point boundary value problems and conducting dynamic collision checks. As a result, the cost of edge evaluation dominates the overall replanning process, especially for motion planning problems in high-dimensional configuration spaces, where the RRG must be dense enough to contain a solution.
The lazy search technique has been introduced to mitigate the problem of excessive edge evaluations [16]. It assumes that a heuristic function can efficiently compute a lower bound on edge weight and uses the lower bound to estimate whether an edge has the potential to improve the solution. This approach defers the evaluation of the actual cost of an edge until it is deemed pertinent to the solution. Various sampling-based algorithms [17, 18, 19] have exploited the lazy search strategy to enhance their planning efficiency. Regrettably, none can achieve both the replanning efficiency of lifelong planning techniques and the edge evaluation efficiency of lazy search techniques.
In this paper, we propose an asymptotically optimal lifelong sampling-based algorithm, Lazy Lifelong Planning Tree (LLPT), for online motion planning in dynamic environments where edge evaluation is expensive. The algorithm integrates the lazy search framework and the lifelong planning framework. Specifically, LLPT approximates the problem domain by an RRG with non-overestimating heuristic edges. The actual edge evaluations of the RRG are restricted to those with the potential to be part of the solution. Whenever the RRG changes (e.g., the actual cost of some edges are evaluated or the RRG is refined), a novel informed rewiring cascade is performed to quickly repair the search tree that spans the RRG to reflect the new information and update the solution. We show that LLPT outperforms several state-of-the-art planning algorithms (RRT[6], Informed RRT[8], BIT[19], RT-RRT[14], RRTX[15]) over a set of motion planning problems, including both static and dynamic ones.
II Related Works
II-A Lifelong Sampling-based Algorithms
Over the past decades, lifelong sampling-based planning for dynamic motion planning problems has evolved through various iterations. Rather than discarding all previous samples like brute-force replanning, some algorithms selectively eliminate invalid search tree branches obstructed by dynamic obstacles [9, 10]. They then incrementally densify the RRG until a new solution is derived. More efficient replanning schemes only remove invalid edges of the spanning tree and preserve the separate branches for further planning. For instance, LRF retains a forest of spanning trees that are dynamically split, regrown, and merged to adapt to obstacles and robot movements [11]. Similarly, MP-RRT retains obstructed branches and actively seeks to reconnect them to the root tree, bridging the gap between the start and goal configurations [12]. RRTFND reduces replanning effort by only reconnecting or regrowing functional disconnected branches of the previous solution path [13].
However, the aforementioned approaches are lack of formal guarantees on solution optimality. Distinguished from previous works, RT-RRT[14] and RRTX[15] were designed to address dynamic shortest-path planning problems (DSPP), offering an asymptotically optimal solution given sufficient replanning time. These methods leverage rewiring cascades to adapt the search tree and navigate around dynamic obstacles. In particular, RT-RRT rewires the entire search tree structure starting from the root, which can be computationally expensive when there are many nodes in the tree. On the other hand, RRTX uses a local rewiring cascade, which focuses only on rewiring the parts of the search tree affected by dynamic changes. This makes it more efficient in scenarios where the environment experiences minor alterations.
It is worth noting that the local rewiring cascade can be enhanced in efficiency by ordering rewiring based on the potential solution quality and constraining rewiring only to the informed node set relevant to the solution, akin to methodologies employed in informed graph-based search techniques (i.e., LPA, D). In the later section, we propose an informed local rewiring cascade and integrate it into the proposed algorithm.
II-B Sampling-Based Planning with Lazy Collision Detection
The lazy sampling-based planning is based on the assumption that a heuristic function can estimate a lower bound on edge costs effectively. The lower bound is used to avoid unnecessary edge evalution. Lazy PRM and Lazy RRG [17] build a roadmap of feasible configurations like PRM and RRG [6], but with the difference that edges are not immediately checked for collision. Only edges that are on a better candidate path to the goal are checked for collision. If an edge is found to collide, it is removed from the roadmap, and the candidate path is updated. Similarly, Lower Bound Tree-RRT (LBT-RRT) [18] proposed the lower bound graph $\mathcal{G}_{lb}$, an RRG with lazily-evaluated edges, to approximate the problem domain. $\mathcal{G}_{lb}$ poses a lower bound invariant to the nodes in the search tree, i.e., the solution cost of the nodes in the search tree is always smaller than their solution cost via $\mathcal{G}_{lb}$. An edge is only evaluated if its lazy insertion into $\mathcal{G}_{lb}$ causes the violation against the lower bound invariant. Batch Informed Tree (BIT) [19] inserts a batch of samples concurrently into the search graph with lazily evaluated edges and prioritizes the evaluation of the new edges with minor lazily-evaluated solution cost. This mechanism boosts evaluation efficiency, albeit at the expense of queuing up edges in the process.
Our work brings together lifelong sampling-based algorithms and lazy search techniques to harvest replanning and edge evaluation efficiency when solving DSPP. Our algorithm outperforms existing lifelong and lazy sampling-based algorithms in terms of planning efficiency for both static and dynamic planning problems.
III Lazy Lifelong Planning Tree
In this section, we first introduce the notations that will be used throughout the paper. In Section III-B, we highlight the informed rewiring cascade, a crucial component of LLPT that efficiently rewires the underlying graph upon detecting modifications. We then present the overall algorithm framework, with a particular focus on the lazy search strategy that is used to delay collision check. Finally, we describe the graph densification strategy used to enhance the solution.
III-A Notations and Problem Definitions
Assume that a robot begins its journey from $v_{start}$ to $v_{goal}$. An optimal path planning problem is to find a path $\pi^{*}(v_{start},v_{goal})$, an ordered set of distinct vertices that connects $v_{start}$ and $v_{goal}$ and minimizes the total cost to traverse this path. LLPT solves this problem by constructing a search graph $\mathcal{G}=(V,E)$ with lazily-evaluated edges within the robot’s configuration space $\mathcal{X}$, where $V$ represents the node set and $E\subseteq V\times V$ denotes the edge set. To clarify, each node $v\in V$ implies the corresponding configuration $x\in\mathcal{X}$ and can be directly input into distance functions. For each node, LLPT stores its neighbor nodes in the memory for future reference. The search tree of $\mathcal{G}$ is denoted by $\mathcal{T}=(V_{\mathcal{T}},E_{\mathcal{T}})$, rooted at $v_{goal}$, where $V_{\mathcal{T}}\subseteq V$ and $E_{\mathcal{T}}\subseteq E$. As mentioned in [15], Rooting $\mathcal{T}$ in $v_{goal}$ allows the same tree and root to be used for the entire navigation, avoiding rewiring the tree from the root while the robot moves.
For each edge $e\in E$ of $\mathcal{G}$, a weight function $w:E\rightarrow(0,\infty]$ assigns a positive real number, including infinity, to this edge, e.g., the distance/time/energy cost for traversing this edge. If traversing the edge is infeasible, the edge weight will be infinity. Denote the cost of traversing the local trajectory of an edge in the free space as $\hat{w}(e)$, and $E_{eval}$ be the set of evaluated edges at the current replanning iteration, the edge weight of $\mathcal{G}$ is defined by a lazy weight function
$\bar{w}(e)=\begin{cases}w(e),&\text{if }e\in E_{eval},\\ \hat{w}(e),&\text{else.}\end{cases}$ (1)
The solution path should minimize the total weight, i.e., $\pi^{*}(v_{start},v_{goal})=\arg\min_{\pi\in\mathcal{G}}\sum_{e\in\Pi}w(e)$, where $\Pi$ is the set of all finite cost paths connecting $v_{start}$ to $v_{goal}$ in $\mathcal{G}$.
1 $E,E_{\mathcal{T}}\leftarrow\phi$; $V\leftarrow\{v_{goal},v_{start}\}$,$V_{\mathcal{T}}\leftarrow\{v_{goal}\}$ ;
2
[…]
4
5while $v_{start}\neq v_{goal}$ do
6       Update the environment information;
7
8      $E_{eval}\leftarrow\phi$, $E_{collided}\leftarrow\phi$;
9
10      Update the edge weight of $\mathcal{G}$ based on Eq.1;
11
12      $E_{modified}\leftarrow$ the set of edges with modified weight;
13
14      for $e=(v,u)\in E_{modified}$ do
15             UpdateNode ($v,u$);
16
17
18      while TimePermits() do
19             repeat
20                   $\pi\leftarrow$ComputeShortestPath ();
21                   if $v_{start},v_{goal}\in\pi$ then
22                         $V_{collided}=$EvaluateEdge($\alpha$, $\pi$);
23                         for $v\in V_{collided}$ do
24                               $lmc(v)\leftarrow\infty$;
25                               RemoveFromTree($v$);
26                               $\mathcal{Q}$.update($v$);
27                               PropagateCostToLeave ($v$);
28
29
30                  else
                         break ;
                          // No solution exist.
31
32
33            until $\pi\subset E_{eval}$ and $\pi\cap E_{collided}=\phi$;
34             ExtendSearchGraph ();
35
36
37      if robot is moving then
38             $v_{last}\leftarrow v_{start}$;
39             Update $v_{start}$;
40             $k_{m}\leftarrow k_{m}+w(v_{last},v_{start})$;
41
42
43
Algorithm 1 Lazy Lifelong Planning Tree
1 $E_{uneval}\leftarrow$ unevaluated edges in $\pi$;
2 $V_{collided}\leftarrow\phi$;
3 if $|E_{uneval}|>\alpha$ then
4       $E_{uneval}\leftarrow$ first $\alpha$ edges in $E_{uneval}$ closest to $v_{goal}$;
5
6for $e(v,u)\in E_{uneval}$ do
7       if $e$ is collided with some obstacles then
8             RemoveFromTree($v$);
9             $lmc(v)\leftarrow\infty$;
10             $E_{collided}\xleftarrow{+}\{e\}$, $V_{collided}\xleftarrow{+}\{v\}$;
11
12      $E_{eval}\xleftarrow{+}\{e\}$;
13
14return $V_{collided}$;
Algorithm 2 EvaluateEdge($\pi$, $\alpha$)
1 $lmc(v)\leftarrow\infty$, $V_{\mathcal{T}}\xleftarrow{-}v$, $\mathcal{Q}$.update($v$);
2 for $u\in\textnormal{{Child(}}\textnormal{\emph{$v$}}\textnormal{{)}}$ do
3       PropagateCostToLeave($u$);
4
[…]
1 while $|\mathcal{Q}|>0$ and $(lmc(v_{start})>g(v_{start})$ or $lmc(v_{start})=g(v_{start})=\infty$ or $v_{start}\in\mathcal{Q}$ or $\mathcal{Q}$.topkey() $<$ CalculateKey($v_{start}$) do
2       $k_{old}\leftarrow\mathcal{Q}$.topkey();
3       $v\leftarrow\mathcal{Q}$.pop();
4       if $k_{old}<$CalculateKey($v$) then
5             $\mathcal{Q}$.update($v$);
6
7      else
8             if $lmc(v)>g(v)$ or $lmc(v)=g(v)=\infty$ then
9                   for $u\in V_{\mathcal{T}}\cap N(v)$ do
10                         if $lmc(v)>\bar{w}(v,u)+lmc(u)$ then
11                               MakeParentOf($v,u$);
12                               $lmc(v)\leftarrow\bar{w}(v,u)+lmc(u)$;
13
14
15
16
17            if $lmc(v)\neq\infty$ then
[…]
19                         if $lmc(u)>\bar{w}(u,v)+lmc(v)$ then
20                               MakeParentOf($u,v$);
21                               $lmc(u)\leftarrow\bar{w}(u,v)+lmc(v)$;
22                               $\mathcal{Q}$.update($u$);
23
24
25
26             $g(v)\leftarrow lmc(v)$;
27
28return path from $v_{start}$ to $v_{goal}$;
Algorithm 4 ComputeShortestPath()
1 if $lmc(v)>\bar{w}(v,u)+lmc(u)$ then
2       MakeParentOf($v$,$u$);
3       $lmc(v)\leftarrow\bar{w}(v,u)+lmc(u)$;
4       $\mathcal{Q}$.update($v$);
5
Algorithm 5 UpdateNode($v$)
1 $v_{rand}\leftarrow$ RandomSample();
2 $v_{nearest}\leftarrow$ Nearest($v_{rand}$);
3 $v_{new}\leftarrow$Steer($v_{nearest},v_{rand}$, $\delta$);
4 if $v_{new}$ is in the free configuration space then
5       $V\xleftarrow{+}\{v_{new}\}$;
6
7for $u\in N(v_{new})$ do
8       if local trajectory exists between $v$ and $u$ then
9             $E\xleftarrow{+}\{e(v_{new},u)\}$;
10             UpdateNode($v,u$);
11
12      if local trajectory exists between $u$ and $v$ then
13             $E\xleftarrow{+}\{e(u,v_{new})\}$;
14
15
16
Algorithm 6 ExtendSearchGraph()
III-B Informed Rewiring Cascade
LLPT dynamically adjusts the solution path by rewiring $\mathcal{T}$ when the edge weights of $\mathcal{G}$ change. Drawing inspiration from RRTX [15] and RRT [7], LLPT employs two cost-to-goal estimates, g($v$) and lmc($v$), for each node to detect local inconsistencies arising from graph modifications. A node $v$ is called consistent if g($v$) = lmc($v$)<absent<$\infty$< ∞, and inconsistent otherwise. An inconsistent node is further categorized as locally overconsistent if lmc($v$) $<$ g($v$), or locally underconsistent if lmc($v$) $>$ g($v$) or lmc($v$) = g($v$) $\infty$. The g-value is the accumulated cost-to-goal by traversing the previous search tree, while the lmc-value is the cost-to-goal based on the up-to-date graph topology and is potentially better informed than the g-value. The lmc-value satisfies the following relationship when all nodes are locally consistent:
$lmc(v)\leftarrow\begin{cases}0,&\text{if }v=v_{goal},\\ \min_{u\in N(v)}\bar{w}(v,u)+lmc(u),&\text{otherwise}.\end{cases}$ (2)
Then the shortest path from $v_{start}$ to $v_{goal}$ can be found by traversing iteratively from $v_{start}$ to the neighbor node $u$ that minimizes $\bar{w}(v,u)+lmc(u)$.
Modifications on edge weights potentially trigger changes in the lmc-values of some nodes and render these nodes inconsistent. LLPT collects and stores the inconsistent nodes in a priority queue $\mathcal{Q}$. The priority of each node in the queue is based on the order of their associated keys. If a node $v$ becomes underconsistent, which happens when it solution path to goal is evaluated to be in collision, LLPT will use the PropagateCostToLeave procedure (Algorithm 3) to propagate the cost ascendance to its leaf nodes. Afterwards, both $v$ and its leaf nodes will be added to the priority queue $\mathcal{Q}$. On the other hand, if a node $v$ becomes overconsistent, which can happen when some of its outgoing edges have decreasing weight or when $\mathcal{G}$ is extended, $v$ will be added to $\mathcal{Q}$.
The ComputeShortestPath procedure, which is provided in Algorithm 4, is called to make the spanning tree consistent again by operating on the inconsistent nodes stored in $\mathcal{Q}$ iteratively. For a underconsistent node, ComputeShortestPath iteratively searches among its neighbor nodes, and the underconsistent node is rewired to a new parent node if the new parent node resulting in a better cost-to-goal (line 4-4, Algorithm 4). For a overconsistent node, its local inconsistency is disseminated to its neighbors. Specifically, if $v$ emerges as a better parent node for a neighbor than its current parent, the neighbor’s parent is updated to $v$, and the neighbor will be included in $\mathcal{Q}$ for further processing until consistency is restored within the nodes relevant to the solution (line 4-4, Algorithm 4).
To save computational cost, we only operate on inconsistent nodes with the potential to be in the solution path. These nodes are called promising in [7]. For a promising node $v$, an admissible cost estimate of the path from $v_{start}$ to $v_{goal}$ constrained to pass through it is smaller than the current best solution cost. Define the function $h(v_{start},v):v\rightarrow\mathbb{R}^{+}$ as an admissible heuristic estimate of the cost from $v_{start}$ to $v$, i.e., $h(v_{start},v)\leq w(v_{start},v)$, and the key of a node as $k(v)=[k_{1}(v),k_{2}(v)]$, where $k_{1}(v)=\min(lmc(v),g(v))+h(v)$ and $k_{2}(v)=\min(lmc(v),g(v))$, the definition of the promising node set $V_{prom}$ is given as follows:
$V_{prom}=\{v\in V|k(v)\dot{<}k(v_{start})\},$ (3)
where $\dot{<}$ is a lexicographical comparator with $k(v)\dot{<}k(u)$ if either $k_{1}(v)<k_{2}(u)$ or $k_{1}(v)=k_{1}(u)\land k_{1}(\mathbf{v})<k_{2}(u)$. The node in $\mathcal{Q}$ with the smallest key is regarded as the most promising one and will be popped out from $\mathcal{Q}$ and processed first until there is not node with smaller key than $v_{start}$ and $v_{start}$ has found a solution path to $v_{goal}$ (line 1, Algorithm 4). Since $v_{start}$ is updated as the robot moves, $h(v_{start},v)$ is changing and thus the key used by the priority queue. To avoid having to reorder the queue, we adopt the key updating strategy as D Lite[5], where a constant $k_{m}$ (initialized as 0) that is added up with the cost of the robot movement at each replanning iteration is used to mitigate the priority inconsistency of the out-of-date priority in $\mathcal{Q}$ with a newly-inserted vertex (line 1, Algorithm 1). The key of a newly-inserted vertex $k(v)$ is calculated by $k_{1}(v)=\min(lmc(v),g(v))+h(v_{start},v)+k_{m}$ and $k_{2}(v)=\min(lmc(v),g(v))$. For each node popped from the top of $\mathcal{Q}$, its key will be recalculated, and the node will be reinserted into $\mathcal{Q}$ with updated priority if its old key is smaller than the new one (line 4-4, Algorithm 4).
III-C Algorithmic framework
The proposed algorithm is provided in Algorithm 1 with its major sub-routines outlined in Algorithms 2-6. Notations such as $A\xleftarrow{+}B$ represents $A\leftarrow A\cup B$, and $A\xleftarrow{-}B$ represents $A\leftarrow A\backslash B$. Set cardinality is denoted by ||$\cdot$|| ⋅ |. $\textnormal{{Child}}(v)$ and $\textnormal{{Parent}}(v)$ retrieve child nodes and the parent of node $v$ in $\mathcal{T}$, respectively. $N(v)$ returns the neighboring nodes of $v$ in $\mathcal{G}$. The priority queue $\mathcal{Q}$ stores nodes sorted in ascending order of their keys. $\mathcal{Q}.\textnormal{{update}}(v)$ insert or re-order node $v$ based on its key value. $\mathcal{Q}.\textnormal{{pop}}()$ pops and returns the vertex with the highest priority in $\mathcal{Q}$. $\mathcal{Q}.\textnormal{{topkey}}$ returns the key of the vertex with the highest priority in $\mathcal{Q}$. The function $\textnormal{{MakeParentOf}}(v,u)$ establishes a child-parent relationship between $v,u$ and adds $v$ to $\mathcal{T}$. $\textnormal{{RemoveFromTree}}(v)$ eliminates the child-parent relationship between $v$ and its parent node and removes $v$ from $\mathcal{T}$. The comparison operator with $\infty$ is defined such that $x<\infty$ is true for any $x\neq\infty$, otherwise false.
LLPT initiates by setting the goal node $v_{goal}$ with lmc($v_{goal}$)=g($v_{goal}$)=0 and the start node $v_{start}$ with lmc($v_{start}$)=g($v_{start}$)$\infty$. The replanning procedure (line 1-1, Algorithm 1) iterates indefinitely at a user-defined frequency until the robot reaches the goal configuration. At the start of each replanning cycle, the search graph $\mathcal{G}$ is refreshed with lazily-computed edges (line 1, Algorithm 1). For edges with updated weights, their starting nodes are rewired using UpdateNode (line 1-1, Algorithm 1). Each replanning cycle comprises two primary loops: the outer loop (line 1-1, Algorithm 1) and the inner loop (line 1-1, Algorithm 1). The inner loop serves as the main search process, rewiring the search tree $\mathcal{T}$ and identifying the shortest collision-free path. This loop alternates between rewiring $\mathcal{T}$ to compute the shortest path and evaluating the edges along the shortest path until a collision-free shortest path is discovered. The procedure EvaluateEdge evaluates the first $\alpha$ unevaluated edges of the shortest path, where $\alpha\in\mathbb{N}^{*}$ is a user-specified parameter to balance the competing computational costs of edge evaluation and rewiring, and returns the collided nodes. For the collided nodes, their lmc-value is increased to infinity. Then, they are removed from $\mathcal{T}$ and inserted into $\mathcal{Q}$ (line 1-1, Algorithm 1). The cost-to-goal increasing information is propagated to their descendants via the procedure PropagateCostToLeave (line 1, Algorithm 1). The inner loop then continues to rewire the search tree and updates the solution (line 1, Algorithm 1). If a solution regarding current $\mathcal{G}$ is not found, the inner loop reports failure and terminates. In case of early termination of the inner loop with remaining time, the search graph $\mathcal{G}$ is extended in the outer loop to enhance potential solution quality (line 1, Algorithm 1).
III-D Search Graph Densification
The ExtendSearchGraph procedure (Algorithm 6) extends the underlying search graph $\mathcal{G}$ to improve both completeness and optimality of the solution. $\mathcal{G}$ is grew to incrementally approximate the problem domain by sampling a random point $v_{rand}$ from $\mathcal{X}$ (line 6, Algorithm 6) and extending some parts of the graph towards $v_{rand}$. First, the closest node in the graph $v_{nearest}$ is found, then a new sample $v_{new}$ is generated by steering $v_{nearest}$ towards a randomly sampled point $v_{rand}$ by distance of $\delta$ (line 6, Algorithm 6). Then, the lazily-evaluated edges among $v_{new}$ and its neighbor nodes are added to $\mathcal{G}$ (line 6-6, Algorithm 6). In this paper, the neighbor nodes are found by searching for nodes that are within radius $r$ of $v_{new}$, which is given in [6] by
$r=\min\left\{2\left(1+\frac{1}{d}\right)^{\frac{1}{d}}\left(\frac{\mu\left(\mathcal{X}_{\text{free }}\right)}{\zeta_{d}}\right)^{\frac{1}{d}}\left(\frac{\log(N)}{N}\right)^{\frac{1}{d}},\delta\right\}$ (4)
where N=$|V|$, $d$ is the dimension of the problem, $\mu\left(\mathcal{X}_{free}\right)$ is the Lebesgue measure of the free configuration space $\mathcal{X}_{free}$, $\zeta_{d}$ is the volume of a unit ball in $\mathbb{R}^{d}$. The lmc- and g-value of each new node are initialized as $\infty$. If $v_{new}$ is able to yield a solution path via one of its neighbor node, its lmc-value will be updated, and the node will be inserted into $\mathcal{Q}$ (line 6, Algorithm 6). Later, the ComputeShortestPath procedure will propagate the new information due to the extension across the graph and search for a better solution for promising nodes.
IV Simulation Results
In this section, we present simulation results comparing LLPT against RRT[6], Informed RRT[8], BIT[19], RRTX[15], and RT-RRT[14] to verify the performance of LLPT. We compare them for static and dynamic path planning problems in $\mathbb{R}^{2}$, $\mathbb{R}^{3}$. The length for each dimension is 30.0. The $\epsilon$ parameter of RRTX is set as 0 for all simulations. BIT samples 150 nodes per batch regardless of problem dimension. The Euclidean norm is used as the weight and the heuristic functions for all problems. The edge resolution is set as 0.02 for collision check. The saturate parameter $\delta$, equivalent to the maximum edge length of the search graph, is set as 2 for $\mathbb{R}^{2}$ and 3 for $\mathbb{R}^{3}$. The simulations were developed in Python 3.0 and ran on a Macbook with an M2 chip and 16GB RAM.
Figure 1: The x-y plane of the obstacle configurations (represented by black regions) of the static planning problems. The red dot and the blue dot denotes the start and the goal configuration, respectively.
IV-A Static Planning Problems
The planners were tested on two abstract problems with different obstacle configurations. The x-y plane of the obstacle configurations is visualized in Fig.1. The start and goal are located at $[10,15,15,...,15]^{T}$ and $[20,15,15,...,15]^{T}$ for the problem illustrated in Fig.1(a), and $[8,8,8,...,8]^{T}$ and $[23,23,23,...,23]^{T}$ for the problem illustrated in Fig.1(b). The evaluation parameter $\alpha$ of LLPT is set as 1. Each planner was tested with 50 different pseudo-random seeds.
(a) $\mathbb{R}^{2}$, Fig.1(a)
(b) $\mathbb{R}^{2}$, Fig.1(b)
(c) $\mathbb{R}^{3}$, Fig.1(a)
(d) $\mathbb{R}^{3}$, Fig.1(b)
[…]
(a) Scene 1
(b) Scene 2
(c) Scene 3
Figure 3: The x-y plane of the obstacle configurations (represented by black regions) of the dynamic planning problems. The red dot and the black dot denotes the start and the goal configuration, respectively.
The simulation results are displayed in Fig.2, showcasing the medium solution cost plotted against run time. LLPT emerges as the top performer among all algorithms, excelling in both the efficiency of initial solution discovery and the convergence rate towards the optimal solution. These results demonstrate the efficacy of lazily delaying edge evaluations in solving motion planning problems. It’s worth noting that LLPT and RRTX outperform informed sampling-based methods (Informed RRT and BIT) in scenarios depicted in Fig.1(a), where the solution path is twisted. In such cases, the informed set comprises a considerable number of nodes, and the informed sampling strategies underperform. However, for the problems depicted in Fig.1(b), BIT displays superior performance over RRTX in $\mathbb{R}^{2}$, and marginally outperforms both RRTX and LLPT in $\mathbb{R}^{3}$ as run time increases. These findings suggest a potential avenue for performance enhancement by integrating LLPT with informed sampling strategies for scenarios with simple obstacle structures.
IV-B Dynamic Planning Problems
We compared the replanning efficiency of LLPT with RRTX and RT-RRT in solving dynamic path planning problems in $\mathbb{R}^{2}$, $\mathbb{R}^{3}$. During planning, the obstacle configurations change from Scene 1 to Scene 3 (Fig.3). The start and goal are located at $[15,1,15,...,15]^{T}$ and $[15,15,15,...,15]^{T}$, respectively. To ensure a fair comparison, all planners assume that the problem domain is free and initialize by generating an RRG $\mathcal{G}$ that approximates the problem domain and surely contains a solution for all cases. Specifically, $\mathcal{G}$ has 3000 for $\mathbb{R}^{2}$ and 9000 nodes for $\mathbb{R}^{3}$. Each planner then undergoes three consecutive planning iterations on $\mathcal{G}$ to adapt their solutions to the changing environment. Note that the environmental change interval is more significant than the time required for replanning to ensure that the planners have enough time to optimize their solutions. Each planner was run 20 times with different pseudo-random seeds.
The average number of edge evaluations, the average number of node expansions, the average time cost, and the average solution cost for initialization and each replanning episode are tabulated in Table I. In this paper, the number of node expansions refers to the total number of nodes that are popped from the priority queue during the replanning procedure. The results demonstrate that LLPT is able to replan successfully in all cases with significantly fewer edge evaluations while maintaining solution quality, showcasing a marked overall efficiency improvement. It’s worth noting that LLPT avoids unnecessary edge evaluations when updating its solution to Scene 1, unlike RRTX and RT-RRT, which perform iterative evaluations across the entire graph even when obstacle movement does not impact the solution.
Moreover, the impact of the evaluation parameter $\alpha$ on LLPT performance is explored in Table 1, revealing its influence on edge evaluations and node expansions. The paper suggests careful selection of $\alpha$ to optimize LLPT performance for specific problems, although the methodology for selecting the ideal value is beyond the paper’s scope.
Figure 4: LLPT planning results to find the shortest path in a dynamic environment. The purple lines are the evaluated edges. The blue lines represent the edges belonging to the search tree. The red line represents the solution path.
TABLE I: The average number of edge evaluations, the average number of vertex expansions, the average replanning time, and the average solution cost for different planners over three consecutive planning in the dynamic environment shown in Fig.3. The absence of a solution was considered a cost of 200.
$\boldsymbol{\mathbb{R}^{2}}$ LLPT RRTX RTRRT
$\alpha$=1 $\alpha$=3 $\alpha$=100
Initialize
# Eval. 10 10 10 119340 119340
# Expan. 799 799 799 11961 5728
Time (s) 47.5 47.5 47.5 67.7 67.0
Cost 14.1 14.1 14.1 14.1 14.1
1st Replan
# Eval. 10 10 10 119340 119340
# Expan. 0 0 0 304 2677
Time (s) 0.001 0.001 0.001 11.5 13.6
Cost 14.1 14.1 14.1 14.1 14.1
2nd Replan
# Eval. 205 341 415 119340 119340
# Expan. 3846 3530 3294 2104 2693
Time (s) 2.1 1.81 1.74 12.9 14.7
Cost 23.3 23.3 23.3 23.3 23.3
3rd Replan
# Eval. 562 852 974 119340 119340
# Expan. 44817 42799 38216 151 2660
Time (s) 14.4 13.5 12.8 12.8 14.5
Cost 51.8 51.8 51.8 200 51.8
Total Time (s) 64.0 62.8 62.0 104.9 109.8
$\boldsymbol{\mathbb{R}^{3}}$
Initialize
# Eval. 6 6 6 303122 303122
# Expan. 1486 1486 1486 20144 541619
Time (s) 430.0 430.0 430.0 516.7 521.4
Cost 14.8 14.8 14.8 14.8 14.8
1st Replan
# Eval. 6 6 6 303122 303122
# Expan. 0 0 0 823 8067
Time (s) 0.001 0.001 0.001 46.0 48.9
Cost 14.8 14.8 14.8 14.8 14.8
2nd Replan
# Eval. 475 798 839 303122 303122
# Expan. 6000 5766 5724 5914 8072
Time (s) 2.2 2.2 2.2 51.9 54.2
Cost 24.8 24.8 24.8 24.8 24.8
3rd Replan
# Eval. 2087 3359 3629 303122 303122
# Expan. 99080 97759 97131 4100 7989
Time (s) 29.6 27.1 27.4 50.1 52.6
Cost 55.2 55.2 55.2 55.2 55.2
Total Time (s) 461.8 459.3 459.6 664.7 677.1
V Conclusions
We introduce LLPT, an asymptotically optimal lazy lifelong sampling-based path planning algorithm. LLPT combines the replanning efficiency of lifelong planning, the search efficiency of heuristic-guided informed algorithms, and the evaluation efficiency of lazy search algorithms. Our simulations show that LLPT outperforms several sampling-based algorithms for both static and dynamic motion planning problems. These results suggest LLPT is well-suited for time-constrained scenarios, particularly for robots in dynamic environments.